. Benzene and toluene form nearly ideal solutions. At 25.0°C the vapor pressure of pure benzene and pure toluene are 95.1 torr and 28.4 torr, respectively. What is the mole fraction of benzene in the vapor in equilibrium with a solution having mole fraction X benzene = 0.40 at 25.0°C?

(a) 0.77

(b) 0.69

(c) 0.60

(d) 0.40

(e) 0.30
Answer #1

Benzene and toluene form nearly ideal solutions. At 25.0°C the vapor pressure of pure benzene and pure toluene are 95.1 torr and 28.4 torr, respectively. What is the mole fraction of benzene in the vapor in equilibrium with a solution having mole fraction X benzene = 0.40 at 25.0°C?

mole fraction of Xbenzene = 0.40

Xtoulene = 0.60

Psolution = PAXA + PBXB

Psolution = Ptoulene x Xtolulene + Pbenzene x Xbenzene

              = 28.4 x 0.60 + 95.1 x 0.40

              = 55.08 torr

mole fraction of benzene in vapor = PAXA / Ptotal

                                                      = 38.04 / 55.08

mole fraction of benzene in vapor = 0.69
Answer #2

Given:

  • Vapor pressure of pure benzene (Pbenzene) = 95.1 torr

  • Vapor pressure of pure toluene (Ptoluene) = 28.4 torr

  • Mole fraction of benzene in the liquid (Xbenzene, liquid) = 0.40

  • Mole fraction of toluene in the liquid (Xtoluene, liquid) = 10.40=0.60

Step 1: Calculate the partial vapor pressures using Raoult's Law :

Pbenzene=Xbenzene, liquid×Pbenzene=0.40×95.1=38.04 torrPtoluene=Xtoluene, liquid×Ptoluene=0.60×28.4=17.04 torr

Step 2: Find the total vapor pressure :

Ptotal=Pbenzene+Ptoluene=38.04+17.04=55.08 torr

Step 3: Calculate the mole fraction of benzene in the vapor phase (Ybenzene) :

Ybenzene=PbenzenePtotal=38.0455.080.69


 Answer is :

                       (b) 0.69  


source: Self
answered by: Harshwardhan kunal